# Krylov Complexity Meets Confinement

### v1

@@ Abstract
Introduction.—Confinement, a hallmark of non-Abelian gauge theories such as quantum chromodynamics (QCD), refers to the phenomenon whereby fundamental excitations—quarks and gluons—cannot be isolated and instead form bound states such as mesons and baryons [1, 2, 3]. While originally formulated in the context of high-energy physics, confinement has also been shown to emerge in certain low-dimensional condensed-matter systems, where elementary excitations experience an effective linear potential that binds them into composite quasiparticles [4, 5, 6, 7, 8, 9]. A number of experiments have explored various aspects of confinement, both in quasi-one-dimensional compounds [10, 11, 12] and in quench experiments with cold-atom systems  [13, 14, 15].
@@
These observations suggest that confinement places stringent constraints on the propagation of information and correlations in quantum many-body systems. Motivated by this, in this Letter we explore whether signatures of confinement can be identified through measures of quantum complexity. In particular, we study Krylov complexity, which quantifies the spread of a time-evolved quantum state or an operator within the Krylov subspace generated by successive applications of the Hamiltonian or a Liovillian on an initial state or an operator [18]. The dynamics of complexity can be mapped to the motion of a quantum particle on a semi-infinite chain, allowing a geometric interpretation of the spread of the wavefunction [19, 20]. Apart from its role in quantifying the spread of quantum information, Krylov complexity, defined for operators or states, has been shown to capture signatures of quantum chaos and thermalization in non-integrable systems [21, 18, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32]. While Krylov operator complexity has been proposed as an order parameter for confinement–deconfinement transitions in large-$N$ gauge theories [33], how the signatures of real-time confinement manifest within the Krylov space of quantum states following a quench remains largely unexplored.
@@
$\hat{H}=-J\sum_{i=1}^{N}\big[\hat{\sigma}^{z}_{j}\hat{\sigma}^{z}_{j+1}+h_{x}\hat{\sigma}^{x}_{j}+h_{z}\hat{\sigma}_{j}^{z}\big].$ (1)
In this work, we set $J=1$ throughout. In the absence of a longitudinal field ($h_{z}=0$), the model maps onto a system of non-interacting fermions and exhibits a quantum phase transition at the critical transverse field strength $h_{x}=1$. For $h_{x}<1$, the system is ferromagnetically ordered, characterized by two degenerate ground states related by a global $\mathbb{Z}_{2}$ symmetry. The low-energy excitations in this regime correspond to freely propagating domain walls (kinks) that separate regions of opposite magnetization and move ballistically in time. Kormos et al. [16] showed that a small longitudinal field $h_{z}$ breaks the $\mathbb{Z}_{2}$ symmetry, lifting the ferromagnetic degeneracy and inducing a linear confining potential that binds domain walls into meson-like excitations. The resulting bound states form a discrete spectrum whose confinement strength grows with $h_{z}$. Following a quench from a ferromagnetic state, it was demonstrated that the magnetization $\langle\hat{\sigma}^{z}(t)\rangle$ exhibits oscillations even for weak fields, while the entanglement entropy rapidly saturates, reflecting confinement dynamics. Interestingly, the oscillations in magnetization and entanglement encode information about the underlying meson spectrum, which manifests as distinct peaks in their power spectra [16, 34].
Figure 2: Krylov complexity $C_{k}(t)$ following a quench from a paramagnetic initial state at $h_{x}=2$ to $h_{x}=1.75$ within the paramagnetic phase, shown for different values of the longitudinal field $h_{z}$. The observed increase in the amplitude of complexity with increasing $h_{z}$ indicates the absence of confinement. The system size used is $L=14$.
@@
We now turn to quenches within the paramagnetic phase ($h_{x}>1$) starting from a paramagnetic state at $h_{x}=2$, as shown in Fig. 2. In sharp contrast to the ferromagnetic regime, the dynamics without a longitudinal field $h_{z}$ displays a markedly suppressed Krylov complexity with only small-amplitude fluctuations, characteristic of free-fermion behavior. Increasing $h_{z}$ leads to a gradual growth in the overall amplitude of the complexity, reflecting enhanced quantum chaotic dynamics arising from interactions. The paramagnetic phase thus remains non-confining and does not host bound meson-like excitations, underscoring the qualitative difference in its dynamical behavior compared to the ferromagnetic case
@@
Lastly, we investigate the growth of complexity for large quenches across the quantum critical point at $h_{x}=1$. Starting from a paramagnetic initial state at $h_{x}=2$, we quench into the ferromagnetic phase to $h_{x}=0.25$. As shown in Fig. 3, the resulting complexity grows to values orders of magnitude larger than those observed for quenches performed within either the paramagnetic or ferromagnetic phases. The large amplitude of the complexity in this case arises because the quench across the critical point excites a broad continuum of modes and generates strong delocalization of the state in Krylov space, reflecting highly non-perturbative dynamics. Unlike intra-phase quenches, the post-quench dynamics in this case exhibits random fluctuations superimposed on the overall growth once the longitudinal field is introduced. Similar to the behavior observed in paramagnetic quenches, the complexity initially increases with the strength of the longitudinal field but eventually shows a distinct turnover, beyond which it begins to decrease. This crossover may signal the emergence of weak confinement, although the signature remains subdued, likely due to the finite propagation velocities of the mesons induced by the large quench amplitude.
@@
Our results thus establish Krylov complexity as a powerful diagnostic of confinement in low-dimensional quantum systems, bridging concepts from quantum information theory with emergent phenomena traditionally studied in high-energy and condensed-matter physics. Several open questions naturally follow. For instance, how does Krylov complexity behave in lattice gauge theories where confinement emerges from gauge constraints [39, 40, 41, 42, 43, 42, 44, 45, 46, 47, 48]? Can similar diagnostics capture other nonperturbative phenomena such as false-vacuum decay [49, 50], string breaking [51, 52], and quantum many-body scarring [53, 54]? Another promising direction is to investigate how confinement affects operator growth in Krylov space, employing finite-temperature algorithms [55] to probe phenomena close to the ground state.
@@ Acknowledgements.
Figure S2: Power spectrum $S_{k}(\omega)$ of the Krylov complexity for a quench within the ferromagnetic phase to $h_{z}=0.1$ and $h_{x}=0.25$, starting from a fully polarized ferromagnetic state. The high-frequency peaks correspond to the four meson masses $m_{1}=3.662$, $m_{2}=4.127$, $m_{3}=4.48$ and $m_{4}=4.769$ obtained from the semiclassical analysis. The spectrum also captures approximately the relative spacing between meson masses such as $m_{12}$ and $m_{13}$. We consider a system of size $L=14$ with Krylov complexity evolved up to time $t=100$ with a time step of $\Delta t=0.1$ for computing the power spectrum using DFT.

### v2

@@ Abstract
Introduction.—Confinement, a hallmark of non-Abelian gauge theories such as quantum chromodynamics (QCD), refers to the phenomenon whereby fundamental excitations—quarks and gluons—cannot be isolated and instead form bound states such as mesons and baryons [1, 2, 3]. While originally formulated in the context of high-energy physics, confinement has also been shown to emerge in certain low-dimensional condensed-matter systems, where elementary excitations experience an effective linear potential that binds them into composite quasiparticles [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. A number of experiments have explored various aspects of confinement, both in quasi-one-dimensional compounds [14, 15, 16] and in quench experiments with cold-atom systems  [17, 18, 19].
@@
These observations suggest that confinement places stringent constraints on the propagation of information and correlations in quantum many-body systems. Motivated by this, in this Letter we explore whether signatures of confinement can be identified through measures of quantum complexity. In particular, we study Krylov complexity, which quantifies the spread of a time-evolved quantum state or an operator within the Krylov subspace generated by successive applications of the Hamiltonian or a Liovillian on an initial state or an operator [22]. The dynamics of complexity can be mapped to the motion of a quantum particle on a semi-infinite chain, allowing a geometric interpretation of the spread of the wavefunction [23, 24]. Apart from its role in quantifying the spread of quantum information, Krylov complexity, defined for operators or states, has been shown to capture signatures of quantum chaos and thermalization in non-integrable closed and open systems [25, 22, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 3, 4, 40, 41, 42, 43, 44, 45, 46]. While Krylov operator complexity has been proposed as an order parameter for confinement–deconfinement transitions in large-$N$ gauge theories [47], how the signatures of real-time confinement manifest within the Krylov space of quantum states following a quench remains largely unexplored.
@@
$\hat{H}=-J\sum_{j=1}^{N}\big[\hat{\sigma}^{z}_{j}\hat{\sigma}^{z}_{j+1}+h_{x}\hat{\sigma}^{x}_{j}+h_{z}\hat{\sigma}_{j}^{z}\big].$ (1)
In this work, we set $J=1$ throughout and use periodic boundary conditions. In the absence of a longitudinal field ($h_{z}=0$), the model maps onto a system of non-interacting fermions and exhibits a quantum phase transition at the critical transverse field strength $h_{x}=1$. For $h_{x}<1$, the system is ferromagnetically ordered, characterized by two degenerate ground states related by a global $\mathbb{Z}_{2}$ symmetry. The low-energy excitations in this regime correspond to freely propagating domain walls (kinks) that separate regions of opposite magnetization and move ballistically in time. Kormos et al. [1] showed that a small longitudinal field $h_{z}$ breaks the $\mathbb{Z}_{2}$ symmetry, lifting the ferromagnetic degeneracy and inducing a linear confining potential that binds domain walls into meson-like excitations. The resulting bound states form a discrete spectrum whose confinement strength grows with $h_{z}$. Following a quench from a ferromagnetic state, it was demonstrated that the magnetization $\langle\hat{\sigma}^{z}(t)\rangle$ exhibits oscillations even for weak fields, while the entanglement entropy rapidly saturates, reflecting confinement dynamics. Interestingly, the oscillations in magnetization and entanglement encode information about the underlying meson spectrum, which manifests as distinct peaks in their power spectra [1, 48].
These measures are useful for detecting confinement, however they primarily capture the spread of spatial correlations rather than how the state itself evolves. In particular, entanglement entropy is sensitive to the choice of bi-partition, highlighting the need for complementary probes such as complexity, which quantifies the global spreading of the quantum state in Hilbert space.
Figure 2: Krylov complexity $C_{k}(t)$ following a quench from a paramagnetic initial state at $h_{x}=2$ to $h_{x}=1.75$ within the paramagnetic phase, shown for different values of the longitudinal field $h_{z}$. The observed increase in the amplitude of complexity with increasing $h_{z}$ indicates the absence of confinement, sharply differing from the quench response within the ferromagnetic phase. The system size used is $L=14$.
@@
We now turn to quenches within the paramagnetic phase ($h_{x}>1$) starting from a paramagnetic state at $h_{x}=2$, as shown in Fig. 2. In sharp contrast to the ferromagnetic regime, the dynamics without a longitudinal field $h_{z}$ displays a markedly suppressed Krylov complexity with only small-amplitude fluctuations, characteristic of free-fermion behavior and the regular oscillations observed at small $h_{z}$ arise from the relatively small quench applied. Increasing $h_{z}$ leads to a rapid growth in the overall amplitude of the complexity, reflecting enhanced quantum chaotic dynamics arising from interactions. The paramagnetic phase thus remains non-confining and does not host bound meson-like excitations, underscoring the qualitative difference in its dynamical behavior compared to the ferromagnetic case
@@
Next, we investigate the growth of complexity for large quenches across the quantum critical point at $h_{x}=1$. Starting from a paramagnetic initial state at $h_{x}=2$, we quench into the ferromagnetic phase to $h_{x}=0.25$. As shown in Fig. 3, the resulting complexity grows to values orders of magnitude larger than those observed for quenches performed within either the paramagnetic or ferromagnetic phases. The large amplitude of the complexity in this case arises because the quench across the critical point excites a broad continuum of modes and generates strong delocalization of the state in Krylov space, reflecting highly non-perturbative dynamics. Unlike intra-phase quenches, the post-quench dynamics in this case exhibits random fluctuations superimposed on the overall growth once the longitudinal field is introduced. Similar to the behavior observed in paramagnetic quenches, the complexity initially increases with the strength of the longitudinal field but eventually shows a distinct turnover, beyond which it begins to decrease. This crossover may signal the emergence of weak confinement, although the signature remains subdued, likely due to the finite propagation velocities of the mesons induced by the large quench amplitude.
Figure 4: Dependence of the long-time–averaged Krylov state complexity $\overline{C_{k}(t)}$ on the confining longitudinal field $h_{z}$ for the three quench protocols discussed in the main text. (a) Quench entirely within the ferromagnetic phase, from a fully ferromagnetic initial state to $h_{x}=0.25$. The inset shows $\overline{C_{k}(t)}$ plotted on a log–log scale, revealing an approximately linear dependence with slope close to 1-1, indicative of the behavior, $\overline{C_{k}(t)}\propto 1/h_{z}$. (b) Quench entirely within the paramagnetic phase, from a paramagnetic initial state at $h_{x}=2$ to $h_{x}=1.75$. The inset displays $\overline{C_{k}(t)}$ on a log–linear scale, where the asymptotically linear behavior suggests an exponential enhancement of the Krylov complexity with increasing longitudinal field $h_{z}$. (c) Quench from a paramagnetic initial state at $h_{x}=2$ to $h_{x}=0.25$, crossing the quantum critical point at $h_{x}=1$. For all the cases, the system size used is L=14.
Lastly, to understand how the Krylov complexity scales with the strength of the confining longitudinal field $h_{z}$, we compute the long-time–averaged Krylov complexity, defined as
$\overline{C_{k}(t)}=\lim_{T\to\infty}\frac{1}{T}\int_{0}^{T}C_{k}(t)\,dt\,.$ (3)
This quantity characterizes the asymptotic growth of the many-body wavefunction in Krylov space by averaging out short-time fluctuations and transient dynamics. We analyze $\overline{C_{k}(t)}$ for all three quench protocols discussed above. The results are shown in Fig. 4. For quenches conducted entirely within the ferromagnetic phase, the data indicate an inverse scaling of the complexity with the longitudinal field, $\overline{C_{k}(t)}\propto h_{z}^{-1}$. In contrast, for quenches within the paramagnetic phase, the behavior is qualitatively different. In this case, the time-averaged Krylov complexity exhibits an exponential enhancement with increasing $h_{z}$, signaling a rapid growth of state complexity driven by the longitudinal field. Finally, for quenches across the quantum critical point from the paramagnetic phase into the ferromagnetic phase where the dependence of $\overline{C_{k}(t)}$ on $h_{z}$ does not reveal a clear scaling form. Nevertheless, we observe that the Krylov complexity initially increases with $h_{z}$, followed by a bending and eventual saturation to a plateau at larger field strengths. This saturation may be attributed to weak confinement effects, which ultimately limit further growth of the state in Krylov space at long times.
@@
Our results thus establish Krylov complexity as a powerful diagnostic of confinement in low-dimensional quantum systems, bridging concepts from quantum information theory with emergent phenomena traditionally studied in high-energy and condensed-matter physics. Several open questions naturally follow. For instance, how does Krylov complexity behave in lattice gauge theories where confinement emerges from gauge constraints [53, 54, 55, 56, 57, 56, 58, 59, 60, 61, 62]? Can similar diagnostics capture other nonperturbative phenomena such as false-vacuum decay [63, 64] and string breaking [65, 66]. Another promising direction is to investigate how confinement affects operator growth in Krylov space, employing finite-temperature algorithms [67] to probe phenomena close to the ground state.
@@ Acknowledgements.
Figure S2: Power spectrum $S_{k}(\omega)$ of the Krylov complexity for a quench within the ferromagnetic phase to $h_{z}=0.1$ and $h_{x}=0.25$, starting from a fully polarized ferromagnetic state. The high-frequency peaks correspond to the four meson masses $m_{1}=3.662$, $m_{2}=4.127$, $m_{3}=4.48$ and $m_{4}=4.769$ obtained from the semiclassical analysis. The spectrum also captures approximately the relative spacing between meson masses such as $m_{12}$ and $m_{13}$. We consider a system of size $L=14$ with Krylov complexity evolved up to time $t=50$ with a time step of $\Delta t=0.1$ for computing the power spectrum using DFT.